OpenStudy (anonymous):

can you explain how to find the end behavior asymptote of 1/x^2+1

OpenStudy (jdoe0001):

the wh0?

OpenStudy (jdoe0001):

what do you mean?

OpenStudy (anonymous):

the worksheet says to find the end B.A. how?

OpenStudy (jdoe0001):

B.A?

OpenStudy (jdoe0001):

ohh the ahemm, heheh

OpenStudy (anonymous):

behavior asymptote

OpenStudy (jdoe0001):

I guess you can always just graph it

OpenStudy (anonymous):

so im assuming its y=0?

OpenStudy (jdoe0001):

well, it has a horizontal asymptote and vertical

OpenStudy (jdoe0001):

the numerator's degree is less than the denominator's, thus the horizontal will be at y=0, yes

OpenStudy (jdoe0001):

hmm well.... wait. a sec... x^2+1 has no real zeros, so,I guess no vertical asympotes

OpenStudy (anonymous):

ok...ive lost all this info since precal...now im taking cal and im forgetting simple things lol

OpenStudy (anonymous):

whats the rule for the exponents being equal

OpenStudy (debbieg):

If degree of num'r = degree of den'r, then the horizontal asymptote is the line y = p/q where p is the leading coefficient of the num'r, q is the leading coefficient of the den'r.

OpenStudy (jdoe0001):

\(\bf \cfrac{an^\square}{bd^\square}\implies \textit{horizontal asymptote} = \cfrac{a}{b}\)

OpenStudy (debbieg):

E.g., in a function like \(y=\dfrac{3x-1}{2x+7}\) the HA is y=3/2. the leading terms "take over" everything else for large values of x, and the x^n cancel out.

OpenStudy (anonymous):

ok..thank you

OpenStudy (debbieg):

You're welcome, and welcome to Open Study! :)

OpenStudy (anonymous):

seems like a great site...so bigger exp for numerator means no ha?

OpenStudy (jdoe0001):

right, if the numerator's degree is bigger, then no HA if it's bigger by "1", then it'd be an "oblique" asymptote

OpenStudy (anonymous):

ok...so can you explain how to find vert asymptotes?

OpenStudy (jdoe0001):

ohh, at the zeros of the denominator

OpenStudy (jdoe0001):

the zeros of the denominator so long they don't make the numerator zero

OpenStudy (anonymous):

ok